Unit 1 Review -

(2.) Find f(g(x)) and g(f(x)) given f(x)= 3x+7 and g(x)= 1/3x - 5

(show step by step)


1
Expert's answer
2021-09-10T11:57:12-0400

"f(g(x))=f(\frac{1}{3x-5})\\
=3(\frac{1}{3x-5})+7\\
=\frac{3+21x-35}{3x-5})\\
=\frac{21x-32}{3x-5})\\
g(f(x))=g(3x+7)\\
=\frac{1}{3(3x+7)-5}\\
=\frac{1}{9x+16}"